Assume that it is 2008. You purchased CSH stock for ​$49 one year ago and it is now selling for ​$57. The company has announced that it plans a ​$8 special dividend. You are considering whether to sell the stock​ now, or wait to receive the dividend and then sell.

a. Assuming 2008 tax​ rates, what​ ex-dividend price of CSH will make you indifferent between selling now and​ waiting?

b. Suppose the capital gains tax rate is 25 % and the dividend tax rate is 38 %​, what​ ex-dividend price would make you indifferent​ now?
Answer #1

Answer to question a]

Assuming that it is 2008 and the year of holding is one year, to answer this question, we need to find out the applicable capital gains tax rate and dividend tax rate in 2008 in United States.

In the year 2008, Dividend Tax Rate was 15% on Ordinary dividend and equal to 0% on Qualified Dividend. Qualified Dividend are those which are paid by a corporation operating in the United States, or by a qualified foreign company and the shares were to have been held for at least 60 days during the 121 days period which starts 60 days prior to the ex-dividend date.

In the year 2008, long term capital gains tax rate was 15% in United States and holding period was one year.

CSH stands for Civitas Social Housing PLC and it is the first real estate investment trust dedicated to investing exclusively into existing portfolios of built social homes in England and Wales.This means its a foreign company. There is no information available whether this is a qualified foreign company or not.

Since the CSH stock was bought 1 year back, the holding period criteria of 1 year is met. This qualifies capital gains, if any, to be taxed under long term capital gains tax rate of 15%.

Dividend as declared by the company is $ 8 and capital gains would have been Market Price - Cost = 57 - 49 = $ 8. Effectively, there is no difference in amount of capital gains and dividend.

Therefore, on the basis of information given in the question and tax rates of 2008 we can draw following conclusions :

1] If CSH is a Qualified Foreign Company, then its dividend should have been Qualified Dividend back then and there should be no tax on the same whereas selling the same before dividend would have exposed investor to tax of 15% on Capital Gains = 57-49 = 8 X 15% = $ 1.20. This means investor receives Capital gains $ 8 - tax $ 1.2 = $ 6.8.

In this case, ex-dividend price is equal to = Cum-Dividend Price Less Dividend Less (Difference in Tax on Capital Gains & Tax on Dividend) = 57 - 8 - (1.2 - 0) = $ 47.8 for being indifferent between selling now & holding for dividend. This was the ex-dividend price of CSH at which he would have been indifferent as his total payback in both cases shall be the same.

Corollary :

When sold as before dividend, his inflow = 57 - capital gains tax 1.2 = $ 55.80 and when sold after dividend = Ex-Dividend Price + Dividend = 47.8 + 8 = $ 55.80. Further, there shall be no capital gains tax on sale after receiving dividend as sale price then is less than cost price..

2]  If CSH is not a Qualified Foreign Company, then its dividend should have been treated as Ordinary Dividend back then and there should be a tax of 15% on the same. Therefore, tax out go shall be equal to 8 X 15% = $ 1.2 and and net inflow after tax on dividend = $ 8 - 1.2 = $ 6.8. The capital gains rate tax too being same, it shall be equal to the same rate as currently calculated on dividend as both dividend and capital gains are equal (i.e. $ 8).

In this case too, ex-dividend price should equal to = Cum-Dividend Price Less Dividend Less (Difference in Tax on Capital Gains & Tax on Dividend) = 57 - 8 - (1.2 - 1.2) = $ 49 for being indifferent between selling now & holding for dividend. This was the ex-dividend price of CSH at which he would have been indifferent as his total payback in both cases would have been same.

This was the case where dividend and capital gains are same and also where tax rates on both are same so cum-dividend and ex-dividend price remaining same, the investor remains indifferent to both scenarios. Further, there is no question of capital gains tax on sale after dividend as there no capital gains as sale price and cost price is same.

Answer to question b]

If the Capital Gains Tax Rate is 25% - The effective tax outgo shall be = capital gains 8 X 25% = $ 2 and net inflow = $ 6 which means he effectively would have received net $ 55 (.e. 57 - 2) after tax on capital gains.

If the Dividend Tax Rate is 38% - The effective tax outgo shall be = dividend 8 X 38% = $ 3.04 and net dividend inflow = $ 4.96.

In this case, Investor must have received $ 55 to be indifferent if he held the shares, received the dividend and then sold the shares at ex-dividend price. This means he shall be subject to dividend tax rate fpr dividend and capital gains tax rate for excess over cost price. Then, his ex-dividend price should have been equal to = Cost Price + After Tax Capital Gains on sell after receipt of dividend = Cost Price + [ Net inflow after Capital Gains Tax if sold before dividend - After Tax Dividend - Cost Price) / (1- Capital Gains Tax Rate) ] = 49 + [ (55 - 4.96 - 49) / (1-25%) ] = 49 + (1.04 / 0.75) = $ 50.39. Ideally, this was the ex-dividend price of CSH at which investor would have been indifferent as his total payback in both cases was to going to be same. In this case, to reiterate, he was subject to capital gains tax @ 25% only if he sells before dividend while if sells after dividend, then he was subject to dividend tax @ 35% plus capital gains tax @ 25% on selling after receiving dividend on excess of ex-dividend price over cost price.

Corollary :

When sold before dividend, his inflow = 57 - capital gains tax 2 = $ 55

And

When sold after dividend = Ex-Dividend Price + After Tax Dividend - Capital Gains Tax = 50.39 + (8 - 3.04) - [(50.39 - 49) X 25%] = 50.39 + 4.96 - 0.35 = $ 55